OpenStudy (anonymous):

A number cube is rolled 120 times. The number 4 comes up 47 times. What is the experimental probability of rolling a 4? What is the theoretical probability of rolling a 4? a. 47/120; 1/30 b. 47/120; 1/6 c. 4/47; 1/6 d. 1/6; 47/120

OpenStudy (anonymous):

experimental probability= number of occurences of that particular event/ number of trials= 47/120 theoretical probability= 1/6 as there are 6 different faces

OpenStudy (anonymous):

Thank you :)